Two masses m1 = 5 kg and m2 = 4.8 kg tied to a string are hanging over a light frictionless pulley. What is the acceleration of the masses when lift free to move? ?

Grade:12

1 Answers

Aman wadhwaniya
6 years ago
m1g-T=m1a eq(1)T-m2g=m2a eq(2)Adding eq 1 and 2m1g-m2g=(m1+m2)a5*9.8-4.8*9.8=(5+4.8)49-47.04=9.8a1.96/9.8=a0.2=a